OpenStudy (anonymous):

How do I solve: (using the substitution method) 2x+3y=-7 x=7-5y

OpenStudy (anonymous):

You substitute your x into the first equation for example: 2(7-5y)+3y=-7

OpenStudy (anonymous):

and you get 14-10y+3y=-7

OpenStudy (anonymous):

then you subtract 14 from both sides of th equal sign 14-10y+3y=-7 -14 -14

OpenStudy (anonymous):

-10y+3y=-7

OpenStudy (anonymous):

then combine like terms so you get -7y=-7

OpenStudy (anonymous):

divide each side by -7

OpenStudy (anonymous):

and y= 1

OpenStudy (anonymous):

Now since you've got your y then you plug it in to the x equation now

OpenStudy (anonymous):

x=7-5(1) x=2(1) x=2

OpenStudy (anonymous):

So, y=1 x=2

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

You're welcome